OpenStudy (anonymous):

Sixteen kg of sulfur are needed to make 49 kg of sulfuric acid. How many kg of sulfur are needed to make 40 kg of sulfuric acid? Round to the nearest kg.

OpenStudy (anonymous):

Should be in math. 16/49 = x/40 x = (16/49) 40 etc